OpenStudy (anonymous):

Solve for x: 3|x − 3| + 2 = 14 No solutions x = −1, x = 8.3 x = 0, x = 7 x = −1, x = 7

OpenStudy (anonymous):

Please Help!

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

i got x = 5 and that isn't one of my options i did: 3|x − 3| + 2 = 14 - 2 -2 3|x − 3| = 12 +3 +3 3|x| = 15 3/3|x| = 15/3 |x| = 5

OpenStudy (anonymous):

no you are suppose to divide by 3

OpenStudy (anonymous):

3|x − 3| = 12 |x − 3| = 12/3 |x − 3| = 4

OpenStudy (anonymous):

Ah. Try looking at it like this: 3*(|x-3|)+2=14 note: absolute value bars act like parentheses (the above is redundant but underlines the point)

OpenStudy (anonymous):

x − 3 = 4 and -x+3=4

OpenStudy (anonymous):

so its x = 7? right

OpenStudy (anonymous):

those are the two equations you need to solve

OpenStudy (anonymous):

yeah that's one of the answers. what's the other one?

OpenStudy (anonymous):

x = -1 or 1?

OpenStudy (anonymous):

-x+3=4 solve that

OpenStudy (amistre64):

just trial and error the 4 options to see what fits ....

OpenStudy (anonymous):

Don't ever do trial an error. A waste of time in tests

OpenStudy (anonymous):

-x + 3 = 4 -3 -3 -x = -1 x = 1?

OpenStudy (amistre64):

well, 3 are trialable :)

OpenStudy (anonymous):

-x + 3 = 4 -3 -3 -x = 1 x = -1?

OpenStudy (amistre64):

trial and error is perfect for tests; take an answer optionand see if it fits, if not; youve narrowed it down

OpenStudy (anonymous):

you got the signs wrong.

OpenStudy (anonymous):

then you want the person to guess?

OpenStudy (amistre64):

no, i want the person to "dbl chk" the answer options

OpenStudy (anonymous):

ohh okay i see where i made the mistake so the answer would be: x = −1, x = 7

OpenStudy (anonymous):

correct "brohoof"

OpenStudy (anonymous):

can you help me with 2 more?

OpenStudy (anonymous):

Ok post another question.

OpenStudy (anonymous):

How many solutions does the equation 2|x − 3| + 4 = 4 have? Zero One Two Infinitely Many

OpenStudy (anonymous):

What if you get a question that doesn't have multiple choice options

OpenStudy (amistre64):

3|x − 3| + 2 = 14 does x=0 work? 3|−3| + 2 = 14 9 + 2 = 14 no does x=7 work? 3|7 − 3| + 2 = 14 12 + 2 = 14 yes does x=-1 work? 3|-1 − 3| + 2 = 14 12 + 2 = 14 yes

OpenStudy (amistre64):

if the question has no "options" then waste the time doing it the long way :)

OpenStudy (anonymous):

\[2|x − 3| + 4 = 4 \\ 2|x − 3| + 4-4 = 4-4\\ 2|x − 3| = 0 \\|x − 3| = 0/2 \\ |x − 3| = 0 \]

OpenStudy (anonymous):

two equations means two what?

OpenStudy (anonymous):

2|x − 3| + 4 = 4 - 4 -4 2|x − 3| = 0 /2 /2 |x-3| = 0 + 3 +3 |x| = 3?

OpenStudy (amistre64):

nice shorthand .... reminds me of someone i know :)

OpenStudy (anonymous):

sp their are two solutions right?

OpenStudy (anonymous):

well the two solutions you get are the same.....

OpenStudy (anonymous):

so I guess it's just one.

OpenStudy (anonymous):

ok well last question?

OpenStudy (anonymous):

huh? me short hand? what are you talking about?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve 3|2x + 2| − 7 = 5 for x.

OpenStudy (anonymous):

\[3|2x + 2| − 7 = 5\\ 3|2x + 2| − 7+7 = 5 +7\\3|2x + 2| = 12\\ |2x + 2| = 12/3\\|2x + 2| = 4\]

OpenStudy (anonymous):

the two equations are then 2x+2=4 and -2x-2=4

OpenStudy (anonymous):

solve each equation for x.

OpenStudy (anonymous):

3|2x + 2| − 7 = 5 +7 +7 3|2x + 2| = 12 /3 /3 |2x + 2| = 4 -2 -2 |2x| = 2 /2 /2 |x| = 1?

OpenStudy (anonymous):

|2x + 2| = 4 from this step that you did you have to split it into two equations.

OpenStudy (anonymous):

You can't just subtract 2 from inside the absolute value

OpenStudy (anonymous):

x = -1 x = 1 right?

OpenStudy (anonymous):

oh then what do i do?

OpenStudy (anonymous):

|2x + 2| = 4 turns into 2x + 2 = 4 and -2x - 2 = 4

OpenStudy (anonymous):

you have two equations now. Solve x for each one

OpenStudy (anonymous):

x = 1 x = -1

OpenStudy (anonymous):

no dude you have to solve for each equation 2x + 2 = 4 and -2x - 2 = 4

OpenStudy (anonymous):

i did! and thats what i get

OpenStudy (anonymous):

x = -3 x = 1

OpenStudy (anonymous):

how about that?

OpenStudy (anonymous):

please help @RyanL.

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

Really? Okay well thank you soo much!(: im glad you actually helped me out and not just gave me the answer(: